Sweeney & Allen, a large marketing firm, adjusts its accounts at the end of each month. The following information is available for the year ending December 31.

  1. A bank loan had been obtained on December 1. Accrued interest on the loan at December 31 amounts to $1,310. No interest expense has yet been recorded.
  2. Depreciation of the firm’s office building is based on an estimated life of 30 years. The building was purchased four years ago for $310,000.
  3. Accrued, but unbilled, revenue during December amounts to $67,000.
  4. On March 1, the firm paid $1,400 to renew a 12-month insurance policy. The entire amount was recorded as Prepaid Insurance.
  5. The firm received $15,000 from King Biscuit Company in advance of developing a six-month marketing campaign. The entire amount was initially recorded as Unearned Revenue. At December 31, $4,300 had actually been earned by the firm.
  6. The company’s policy is to pay its employees every Friday. Since December 31 fell on a Wednesday, there was an accrued liability for salaries amounting to $1,900.


a. Record the necessary adjusting journal entries on December 31.

b. By how much did Sweeney & Allen’s net income increase or decrease as a result of the adjusting entries performed in part a? (Ignore income taxes.)
Answer #1

Answer:

a) JOURNAL ENTRIES FOR ADJUSTMENT:
1 Interest expense 1,310
Interest payable 1,310
2 Depreciation expense-Building =310000/(30*12) = 861
Accumulated depreciation-Building 861
3 Accounts receivable 67,000.
Sales revenue 67,000.
4 Insurance expense (1400/12) 116
Prepaid insurance 116
5 Unearned revenue 4,300
Sales revenue 4,300
6 Wages and salaries expense 1,900
Wages and salaries payable

1,900.

b)

Increase in revenue = 67000+4,300= $71,300

Increase in expenses = 1,310+861 +116+1,900 = 4187
Increase in net income = 71,300 - 4187 = $67,113